# DEFT: Differentiable Branched Discrete Elastic Rods for Modeling Furcated DLOs in Real-Time

### v1

@@ IV-E2 Enforcing Attachment at the Junction
IV-E3 Edge Orientation Constraints at the Junction
Many practical wire-harnesses feature junctions that are nearly rigid. This near-rigidity arises because manufacturers often use rigid plastics or other stiff materials at these connection points [37]. This results in minimal rotational or bending freedom. Consequently, the dynamics between the parent and child branches propagate through the junction in an almost rigid manner. To capture this behavior, we model each edge as a rigid body and define a constraint function that computes the relative orientation of each body. We then require that this constraint function is approximately satisfied where ap By applying the following theorem, whose proof can be found in Appendix -D, one can compute $\Delta\hat{\mathbf{x}}^{p,i},\Delta\hat{\mathbf{x}}^{c,1}$ explicitly:
@@ IV-E3 Edge Orientation Constraints at the Junction
$C^{p,i}_{O}(\hat{\mathbf{X}},\Delta\hat{\mathbf{X}})=\|\big{(}\hat{\bm{\Omega}}^{p,i}(\hat{\mathbf{X}},\Delta\hat{\mathbf{X}})+\hat{\bm{\Omega}}_{\Delta}^{p,i}(\hat{\mathbf{X}},\Delta\hat{\mathbf{X}})\big{)}-\\ \big{(}\hat{\bm{\Omega}}^{c,i}(\hat{\mathbf{X}},\Delta\hat{\mathbf{X}})+\hat{\bm{\Omega}}_{\Delta}^{c,i}(\hat{\mathbf{X}},\Delta\hat{\mathbf{X}}\big{)}\|_{2}-\epsilon,$ (25)
where $\hat{\bm{\Omega}}_{\Delta}$ and $\hat{\bm{\Omega}}$ are defined in Appendix -B and $\epsilon>0$ bounds the mismatch between the updated parent–child orientation and the original orientation, ensuring that each junction remains almost rigid. The orientation correction at the junction is obtained by solving the following optimization problem:
@@
When $\Delta_{t}$ is sufficiently small, then one can compute explicit formulas for $\Delta\hat{\mathbf{X}}$ which can be found in Appendix -D.
Note that the junction orientation constraint is satisfied at vertex $p,i$ when $C^{p,i}_{O}(\hat{\mathbf{X}},\mathbf{0})^{2}=0$. For convenience, let the function that computes the minimizer of optimization problem (21) using $\hat{\mathbf{X}}$ be denoted by $D^{p,i}_{O}$ (i.e., $D^{p,i}_{O}:\hat{\mathbf{X}}\mapsto\Delta\hat{\mathbf{X}}$).
IV-E4 Algorithm to Implement Corrections Due To Constraint Enforcement
@@
IV-F Parallel Programming
We implement DEFT using PyTorch [26], taking advantage of its batch operations to execute Algorithm 1. PyTorch’s support for batch-wise computations enables parallel processing of multiple parent and child branch instances, significantly enhancing computational efficiency as we illustrate empirically in Section V-D. Specifically, the first batch is dedicated to executing the parent branch’s algorithm, while the remaining batches handle executions for the children branches.
One challenge arises when the number of nodes in the parent branch ($n^{p}$) differs from the number in a child branch ($n^{c}$). This dimensional mismatch prevents the direct construction of uniform matrices for batch operations. To overcome this, we assume $n^{p}>n^{c}$ and introduce padded matrices and vertices by appending zeros to the child branch’s structure to match the parent branch’s dimensions. This strategy ensures uniform matrix construction across the batches, which facilitates seamless batch-wise operations. During parameter updates, we mask out the gradients from padded elements to prevent them from influencing the learned parameters.
V Experiments and Results
This section presents the real-world evaluation of DEFT, covering: 1) A quantitative accuracy comparison with state-of-the-art methods, including scenarios where grasping occurs at various points along the BDLO, such as parent-branch ends and midpoints; 2) An assessment of DEFT’s inference speed; 3) An ablation study exploring parameter auto-tuning, residual learning, and a naive multi-branch approach that connects multiple single-DLO models; 4) Shape-matching manipulation demonstrations, highlighting DEFT’s real-world manipulation efficacy; 5) Thread insertion manipulation demonstrations, showcasing DEFT’s versatility. Code and data can be found at https://github.com/roahmlab/DEFT. To the best of our knowledge, this will be the first publicly available dataset and code for modeling BDLOs.
V-A Hardware Setup

### v5

@@ B.4.2 Enforcing Attachment at the Junction
$\Delta\hat{\mathbf{x}}^{p,i}=\mathbf{M}^{c,1}(\mathbf{M}^{p,i}+\mathbf{M}^{c,1})^{-1}C^{p,i}_{A}(\hat{\mathbf{X}},\mathbf{0})\frac{(\hat{\mathbf{x}}^{c,1}-\hat{\mathbf{x}}^{p,i})}{||\hat{\mathbf{x}}^{c,1}-\hat{\mathbf{x}}^{p,i}||_{2}},$ (24)
$\Delta\hat{\mathbf{x}}^{c,1}=\mathbf{M}^{p,i}(\mathbf{M}^{p,i}+\mathbf{M}^{c,1})^{-1}\ C^{p,i}_{A}(\hat{\mathbf{X}},\mathbf{0})\frac{(\hat{\mathbf{x}}^{p,i}-\hat{\mathbf{x}}^{c,1})}{||\hat{\mathbf{x}}^{c,1}-\hat{\mathbf{x}}^{p,i}||_{2}}.$ (25)
Proof. To prove Theorem 4, we first formulate the optimization problem:
$\underset{\Delta\hat{\mathbf{X}}}{\min}\quad$ $\frac{1}{2}\left(C^{p,i}_{A}(\hat{\mathbf{X}},\Delta\hat{\mathbf{X}})\right)^{2}$ (26)
s.t. $\Delta\hat{\mathbf{x}}^{j}=\mathbf{0},\quad\forall j\notin\{(p,i),(c,1)\},$ (27)
$\mathbf{M}^{p,i}\,\Delta\hat{\mathbf{x}}^{p,i}+\mathbf{M}^{c,1}\,\Delta\hat{\mathbf{x}}^{c,1}=\mathbf{0},$ (28)
$\mathbf{I}^{p,i}\,\hat{\bm{\Omega}}^{p,i}_{\Delta}(\hat{\mathbf{X}},\Delta\hat{\mathbf{X}})=\mathbf{0}.$ (29)
To solve the above optimization, we introduce Lagrange multipliers $\bm{\lambda}_{\text{l}}\in\mathbb{R}^{3}$ and $\bm{\lambda}_{\text{r}}\in\mathbb{R}^{3}$ associated with constraints (28) and (29), respectively. The corresponding Lagrangian $\mathcal{L}$ can be found as follow:
$\begin{split}\mathcal{L}(\Delta\hat{\textbf{x}}^{p,i},\Delta&\hat{\mathbf{x}}^{c,1},\bm{\lambda}_{\text{l}},\bm{\lambda}_{\text{r}})=\frac{1}{2}\left(C^{p,i}_{A}(\hat{\mathbf{X}},\Delta\hat{\mathbf{X}})\right)^{2}-\\ &\bm{\lambda}^{T}_{\text{l}}(\mathbf{M}^{i}\Delta\hat{\textbf{x}}^{p,i}+\mathbf{M}^{c,1}\Delta\hat{\mathbf{x}}^{c,1})-\bm{\lambda}^{T}_{\text{r}}\mathbf{I}^{p,i}\hat{\bm{\Omega}}^{p,i}_{\Delta}(\hat{\mathbf{X}},\Delta\hat{\mathbf{X}})\end{split}$ (30)
Next, we take the partial derivatives of (30) with respect to $\Delta\hat{\textbf{x}}^{p,i},\Delta\hat{\mathbf{x}}^{c,1},\bm{\lambda}_{\text{l}},\bm{\lambda}_{\text{r}}$. Setting each derivative to zero yields the system of equations (31), (32), (33), and (34).
$\begin{split}&\frac{\partial\mathcal{L}(\Delta\hat{\textbf{x}}^{p,i},\Delta\hat{\mathbf{x}}^{c,1},\bm{\lambda}_{\text{l}},\bm{\lambda}_{\text{r}})}{\partial\Delta\hat{\mathbf{x}}^{p,i}}=\\ &C^{p,i}_{A}(\hat{\mathbf{X}},\Delta\hat{\mathbf{X}})\frac{(\hat{\textbf{x}}^{p,i}+\Delta\hat{\textbf{x}}^{p,i})-(\hat{\mathbf{x}}^{c,1}+\Delta\hat{\mathbf{x}}^{c,1})}{||(\hat{\textbf{x}}^{p,i}+\Delta\hat{\textbf{x}}^{p,i})-(\hat{\mathbf{x}}^{c,1}+\Delta\hat{\mathbf{x}}^{c,1})||_{2}}-\bm{\lambda}^{T}_{\text{l}}\mathbf{M}^{p,i}-\frac{\partial\bm{\lambda}^{T}_{\text{r}}\mathbf{I}^{p,i}\hat{\bm{\Omega}}^{p,i}_{\Delta}(\hat{\mathbf{X}},\Delta\hat{\mathbf{X}})}{\partial\Delta\hat{\mathbf{x}}^{p,i}}=\textbf{0}\end{split}$ (31)
$\begin{split}&\frac{\partial\mathcal{L}(\Delta\hat{\textbf{x}}^{p,i},\Delta\hat{\mathbf{x}}^{c,1},\bm{\lambda}_{\text{l}},\bm{\lambda}_{\text{r}})}{\partial\Delta\hat{\mathbf{x}}^{c,1}}=\\ &C^{p,i}_{A}(\hat{\mathbf{X}},\Delta\hat{\mathbf{X}})\frac{(\hat{\mathbf{x}}^{c,1}+\Delta\hat{\mathbf{x}}^{c,1})-(\hat{\textbf{x}}^{p,i}+\Delta\hat{\textbf{x}}^{p,i})}{||(\hat{\textbf{x}}^{p,i}+\Delta\hat{\textbf{x}}^{p,i})-(\hat{\mathbf{x}}^{c,1}+\Delta\hat{\mathbf{x}}^{c,1})||_{2}}-\bm{\lambda}^{T}_{\text{l}}\mathbf{M}^{c,1}-\frac{\partial\bm{\lambda}^{T}_{\text{r}}\mathbf{I}^{p,i}\hat{\bm{\Omega}}^{p,i}_{\Delta}(\hat{\mathbf{X}},\Delta\hat{\mathbf{X}})}{\partial\Delta\hat{\mathbf{x}}^{c,1}}=\textbf{0}\end{split}$ (32)
$\frac{\partial\mathcal{L}(\Delta\hat{\textbf{x}}^{p,i},\Delta\hat{\mathbf{x}}^{c,1},\bm{\lambda}_{\text{l}},\bm{\lambda}_{\text{r}})}{\partial\bm{\lambda}_{\text{l}}}=-(\mathbf{M}^{i}\Delta\hat{\textbf{x}}^{p,i}+\mathbf{M}^{c,1}\Delta\hat{\mathbf{x}}^{c,1})=\textbf{0}$ (33)
$\begin{split}&\frac{\partial\mathcal{L}(\Delta\hat{\textbf{x}}^{p,i},\Delta\hat{\mathbf{x}}^{c,1},\bm{\lambda}_{\text{l}},\bm{\lambda}_{\text{r}})}{\partial\bm{\lambda}_{\text{r}}}=-\textit{{I}}^{i}\bm{\lambda}^{T}_{\text{r}}\mathbf{I}^{p,i}\hat{\bm{\Omega}}^{p,i}_{\Delta}(\hat{\mathbf{X}},\Delta\hat{\mathbf{X}})=\textbf{0}\end{split}$ (34)
We first observe that solving (33) leads to the following equation:
$\Delta\hat{\textbf{x}}^{c,1}=-(\mathbf{M}^{c,1})^{-1}\mathbf{M}^{p,i}\Delta\hat{\mathbf{x}}^{p,i}$ (35)
This indicates that $\Delta\hat{\textbf{x}}^{p,i},\Delta\hat{\mathbf{x}}^{c,1}$ are colinear. Based on the definition of $\hat{\bm{\Omega}}^{i}_{\Delta}(\hat{\mathbf{X}},\Delta\hat{\mathbf{X}})$ in Appendix B.2, one can see that if $\Delta\hat{\textbf{x}}^{p,i},\Delta\hat{\mathbf{x}}^{c,1}$ are colinear, then $\hat{\bm{\Omega}}^{p,i}_{\Delta}(\hat{\mathbf{X}},\Delta\hat{\mathbf{X}})=\textbf{0}$. Note that this is true regardless of the size of $\Delta_{t}$. Hence, it naturally satisfies (34). Next, summing (31) with (32) yields the following:
$\bm{\lambda}_{\text{l}}\mathbf{M}^{p,i}+\bm{\lambda}_{\text{l}}\mathbf{M}^{c,1}=\textbf{0}$ (36)
As mass matrices are always positive definite. This indicates $\bm{\lambda_{l}}=\textbf{0}$. Based on (31), (35) and (36), solving for $\Delta\hat{\mathbf{x}}^{p,i}$ is equivalent to solving the following equation:
$C^{p,i}_{A}(\hat{\mathbf{X}},\Delta\hat{\mathbf{X}})=\textbf{0}$ (37)
After applying some algebraic manipulations, $\Delta\hat{\mathbf{x}}^{p,i}$ can be found to be:
@@
B.4.3 Edge Orientation Constraints
Many practical wire-harnesses feature junctions that are nearly rigid. This near-rigidity arises because manufacturers often use rigid plastics or other stiff materials at these connection points [39]. This results in minimal rotational or bending freedom. Consequently, the dynamics between the parent and child branches propagate through the junction in an almost rigid manner. To capture this behavior, we model each edge as a rigid body and define a constraint function that computes the relative orientation of each body. We then require that this constraint function is approximately satisfied. One can compute $\Delta\hat{\mathbf{x}}^{p,i},\Delta\hat{\mathbf{x}}^{c,1}$ explicitly:
@@ B.4.3 Edge Orientation Constraints
$C^{p,i}_{O}(\hat{\mathbf{X}},\Delta\hat{\mathbf{X}})=\|\big{(}\hat{\bm{\Omega}}^{p,i}(\hat{\mathbf{X}},\Delta\hat{\mathbf{X}})+\hat{\bm{\Omega}}_{\Delta}^{p,i}(\hat{\mathbf{X}},\Delta\hat{\mathbf{X}})\big{)}-\big{(}\hat{\bm{\Omega}}^{c,i}(\hat{\mathbf{X}},\Delta\hat{\mathbf{X}})+\hat{\bm{\Omega}}_{\Delta}^{c,i}(\hat{\mathbf{X}},\Delta\hat{\mathbf{X}}\big{)}\|_{2}-\epsilon,$ (38)
where $\hat{\bm{\Omega}}_{\Delta}$ and $\hat{\bm{\Omega}}$ are defined in Appendix B.2 and $\epsilon>0$ bounds the mismatch between the updated parent–child orientation and the original orientation, ensuring that each junction remains almost rigid. The orientation correction at the junction is obtained by solving the following optimization problem:
s.t. $\hskip 8.5359pt\mathbf{I}^{p,i}\hat{\bm{\Omega}}_{\Delta}^{p,i}+\mathbf{I}^{c,1}\hat{\bm{\Omega}}_{\Delta}^{c,1}=\mathbf{0}$ (40)
When $\Delta_{t}$ is sufficiently small, explicit formulas for $\Delta\hat{\mathbf{X}}$ can be computed as follows:
$\hat{\bm{\Omega}}_{\Delta}^{p,i}\;=\;\mathbf{I}^{c,1}\,\bigl{(}\mathbf{I}^{c,1}+\mathbf{I}^{p,i}\bigr{)}^{-1}\,C^{p,i}_{O}(\hat{\mathbf{X}},\mathbf{0})\,\frac{\hat{\bm{\Omega}}^{c,1}-\hat{\bm{\Omega}}^{p,i}}{\|\hat{\bm{\Omega}}^{c,1}-\hat{\bm{\Omega}}^{p,i}\|_{2}},$ (41)
$\hat{\bm{\Omega}}_{\Delta}^{c,1}\;=\;\mathbf{I}^{p,i}\,\bigl{(}\mathbf{I}^{c,1}+\mathbf{I}^{p,i}\bigr{)}^{-1}\,C^{p,i}_{O}(\hat{\mathbf{X}},\mathbf{0})\,\frac{\hat{\bm{\Omega}}^{p,i}-\hat{\bm{\Omega}}^{c,1}}{\|\hat{\bm{\Omega}}^{c,1}-\hat{\bm{\Omega}}^{p,i}\|_{2}}.$ (42)
Once $\hat{\bm{\Omega}}_{\Delta}^{p,i},\hat{\bm{\Omega}}_{\Delta}^{c,1}$ are obtained, an approach described in the following are used to update $\hat{\mathbf{x}}^{p,i}$ and $\hat{\mathbf{x}}^{c,1}$ accordingly. Note that the junction orientation constraint is satisfied at vertex $p,i$ when $C^{p,i}_{O}(\hat{\mathbf{X}},\mathbf{0})^{2}=0$. For convenience, let the function that computes the minimizer of optimization problem (20) using $\hat{\mathbf{X}}$ be denoted by $D^{p,i}_{O}$ (i.e., $D^{p,i}_{O}:\hat{\mathbf{X}}\mapsto\Delta\hat{\mathbf{X}}$).
Proof.
Figure 8: An illustration of applying $\hat{\bm{\Omega}}_{\Delta}^{i}$ to rotate the $i$-th segment.
For convenience in the proof, we suppress the arguments to functions for convenience. To prove Theorem 5, we begin by describing how $\hat{\bm{\Omega}}^{i}_{\Delta}$ is applied to rotate the $i$-th segment. Note that $\Delta_{t}$ should be sufficiently small for $\hat{\bm{\Omega}}^{i}_{\Delta}$ to accurately approximate the change in orientation. As shown in Figure 8, the rotation axis coincides with the center of the $i$-th segment. In practice, the rotation of the $i$-th segment is achieved by converting $\hat{\bm{\Omega}}^{i}_{\Delta}$ into the rotation matrix $\hat{\bm{\Omega}}^{i}_{\Delta}\mapsto\mathbf{R}^{i}_{\Delta}$. Consequently, we obtain $\Delta\hat{\mathbf{x}}^{i}$ and $\Delta\hat{\mathbf{x}}^{i+1}$ via
$\Delta\hat{\mathbf{x}}^{i}\;=\;\mathbf{R}^{i}_{\Delta}\,\frac{\bigl{(}\hat{\mathbf{x}}^{i+1}-\hat{\mathbf{x}}^{i}\bigr{)}}{2},$ (43)
$\Delta\hat{\mathbf{x}}^{i+1}\;=\;\mathbf{R}^{i}_{\Delta}\,\frac{\bigl{(}\hat{\mathbf{x}}^{i}-\hat{\mathbf{x}}^{i+1}\bigr{)}}{2}.$ (44)
Because each segment rotates only about its own center, linear momentum is naturally conserved. Thus, we can simplify the optimization problem and write the following optimization problem:
@@
where (38), (39), and (40) are analogous to (23), (26), and (28), respectively. Next, we introduce a Lagrange multiplier and follow the same derivation steps as in (31), (35), and (36), which leads to:
$\hat{\bm{\Omega}}_{\Delta}^{p,i}\;=\;\mathbf{I}^{c,1}\,\bigl{(}\mathbf{I}^{c,1}+\mathbf{I}^{p,i}\bigr{)}^{-1}\,C^{p,i}_{O}(\hat{\mathbf{X}},\mathbf{0})\,\frac{\hat{\bm{\Omega}}^{c,1}-\hat{\bm{\Omega}}^{p,i}}{\|\hat{\bm{\Omega}}^{c,1}-\hat{\bm{\Omega}}^{p,i}\|_{2}},$
$\hat{\bm{\Omega}}_{\Delta}^{c,1}\;=\;\mathbf{I}^{p,i}\,\bigl{(}\mathbf{I}^{c,1}+\mathbf{I}^{p,i}\bigr{)}^{-1}\,C^{p,i}_{O}(\hat{\mathbf{X}},\mathbf{0})\,\frac{\hat{\bm{\Omega}}^{p,i}-\hat{\bm{\Omega}}^{c,1}}{\|\hat{\bm{\Omega}}^{c,1}-\hat{\bm{\Omega}}^{p,i}\|_{2}}.$
Once $\hat{\bm{\Omega}}_{\Delta}^{p,i},\hat{\bm{\Omega}}_{\Delta}^{c,1}$ are obtained, (43) and (44) are used to update $\hat{\mathbf{x}}^{p,i}$ and $\hat{\mathbf{x}}^{c,1}$ accordingly.
B.4.4 Algorithm to Implement Corrections Due To Constraint Enforcement
@@
B.5 Proof of Theorem 2
To prove the theorem, we begin by introducing the curvature binormal and material curvature.
Material Curvatures: The curvature binormal $\kappa\mathbf{b}^{i}$ is traditionally used to represent the turning angle and axis between two consecutive edges:
$\kappa\mathbf{b}^{i}=\frac{2\,\mathbf{e}^{i-1}\times\mathbf{e}^{i}}{\|\mathbf{e}^{i-1}\|_{2}\|\mathbf{e}^{i}\|_{2}+\mathbf{e}^{i-1}\cdot\mathbf{e}^{i}},$ (46)
where $\mathbf{e}^{i-1}$ and $\mathbf{e}^{i}$ are consecutive edge vectors, $\times$ denotes the cross product, and $\cdot$ denotes the dot product. By incorporating the material frame (12) and (13), the curvature binormal (46) is projected onto the material frame to quantify the extent to which the curvature aligns with the frame’s orientation. This projection, which we call the material curvature, provides additional insights into the deformation characteristics, allowing us to distinguish between bending and twisting behaviors within the material frame and is defined as:
$\bm{\omega}^{(i,j)}=\left(\kappa\mathbf{b}^{i}\cdot\mathbf{m}_{1}^{j},\kappa\mathbf{b}^{i}\cdot\mathbf{m}_{2}^{j}\right)^{T}\quad\text{for}\quad j\in\{i-1,i\}.$ (47)
Potential Energy: The potential energy is composed of the bending energy and twisting energy:
$P(\Gamma(\mathbf{X}_{t},\bm{\theta}_{t}),\bm{\alpha}\bigr{)}=P_{\text{bend}}(\Gamma(\mathbf{X}_{t},\bm{\theta}_{t}),\bm{\alpha}\bigr{)}+P_{\text{twist}}(\Gamma(\mathbf{X}_{t},\bm{\theta}_{t}),\bm{\alpha}\bigr{)},$ (48)
where
$\begin{split}&P_{\text{bend}}(\Gamma(\mathbf{X}_{t},\bm{\theta}_{t}),\bm{\alpha}\bigr{)}=\\ &\sum_{i=1}^{n-1}\sum_{j=i-1}^{i}\frac{1}{2}\left(\bm{\omega}^{(i,j)}-\overline{\bm{\omega}}^{(i,j)}\right)^{T}\mathbf{B}^{j}\left(\bm{\omega}^{(i,j)}-\overline{\bm{\omega}}^{(i,j)}\right),\end{split}$ (49)
and
$P_{\text{twist}}(\Gamma(\mathbf{X}_{t},\bm{\theta}_{t}),\bm{\alpha}\bigr{)}=\sum_{i=1}^{n-1}\frac{1}{2}\beta^{i}\left(\theta^{i}-\theta^{i-1}\right)^{2},$ (50)
where $\overline{\bm{\omega}}^{(i,j)}$ denotes the undeformed material curvature, which is calculated when the DLO is in a static state without any external or internal forces applied, and $\mathbf{B}$ and $\beta$, are components of $\bm{\alpha}$, representing the bending stiffness and twisting stiffness, respectively.
Gradient of Potential Energy: With (49) and (50), the gradient of (48) can be derived as following:
$\frac{\partial P(\Gamma(\mathbf{X}_{t},\bm{\theta}_{t}),\bm{\alpha}\bigr{)}}{\partial\theta^{i}}=\frac{\partial P_{bend}(\Gamma(\mathbf{X}_{t},\bm{\theta}_{t}),\bm{\alpha}\bigr{)}}{\partial\theta^{i}}+\frac{\partial P_{twist}(\Gamma(\mathbf{X}_{t},\bm{\theta}_{t}),\bm{\alpha}\bigr{)}}{\partial\theta^{i}}$ (51)
We begin by deriving the first term. Note that $\theta^{i}$ is only relevant in $\bm{\omega}^{(i,i)}$ and $\bm{\omega}^{(i+1,i)}$. Using the chain rule, we obtain:
$\frac{\partial P_{bend}(\Gamma(\mathbf{X}_{t},\bm{\theta}_{t}),\bm{\alpha}\bigr{)}}{\partial\theta^{i}}=\sum_{k=i}^{i+1}\frac{\partial P_{bend}(\Gamma(\mathbf{X}_{t},\bm{\theta}_{t}),\bm{\alpha}\bigr{)}}{\partial\bm{\omega}^{(k,i)}}\frac{\partial\bm{\omega}^{(k,i)}}{\partial\theta^{i}}.$ (52)
To compute $\frac{\partial\bm{\omega}^{(k,i)}}{\partial\theta^{i}}$, we use the identities $\frac{\partial\mathbf{m}_{1}^{i}}{\partial\theta^{i}}=\mathbf{m}_{2}^{i}$ and $\frac{\partial\mathbf{m}_{2}^{i}}{\partial\theta^{i}}=-\mathbf{m}_{1}^{i}$. These lead to:
$\frac{\partial\bm{\omega}^{(k,i)}}{\partial\theta^{i}}=\begin{bmatrix}0&1\\ -1&0\end{bmatrix}\bm{\omega}^{(k,i)}$ (53)
Substituting this result back, we obtain:
$\frac{\partial P_{\mathrm{bend}}(\Gamma(\mathbf{X}_{t},\bm{\theta}_{t}),\bm{\alpha}\bigr{)}}{\partial\theta^{i}}\;=\;\sum_{k=i}^{\,i+1}\bigl{(}\mathbf{B}^{k}\,(\bm{\omega}^{(k,i)}-\bar{\bm{\omega}}^{(k,i)})\bigr{)}^{T}\cdot\begin{bmatrix}0&1\\[6.0pt] -1&0\end{bmatrix}\bm{\omega}^{(k,i)},$ (54)
The second term of (51) can be derived, resulting in:
$\frac{\partial P_{twist}(\Gamma(\mathbf{X}_{t},\bm{\theta}_{t}),\bm{\alpha}\bigr{)}}{\partial\theta^{i}}=\beta^{i}(\theta^{i}-\theta^{i-1})-\beta^{i+1}(\theta^{i+1}-\theta^{i})$ (55)
By substituting (54) and (55) into (51), we obtain the analytical gradient of the potential energy.
Appendix C Supplementary Material for Experiments
C.1 Hardware Setup
Four distinct BDLOs were constructed to validate DEFT’s modeling accuracy as shown at the bottom of Figure 9. The material stiffness description of each BDLO is outlined in Table 4. Note that we introduced stiff child branches attached to the parent branch in BDLO3 and BDLO4. These branches represent the rigid components (e.g., fixations) commonly used in BDLO construction [39]. These attachments help ensure secure mounting, preserve structural integrity, and manage loads in real‐world applications. Note that we selected two child branches for each BDLO to evaluate DEFT’s performance across combinations of branches with varying stiffness. However, our methodology is theoretically scalable to BDLOs with arbitrary tree-like topologies. Spherical markers were attached to each BDLO to facilitate dataset creation for both training and evaluation, with the number of vertices set equal to the number of MoCap markers. Because the markers are present during both training and evaluation, their impact on the BDLO dynamics is inherently accounted within all subsequent analyses. If one wishes to obtain a BDLO dataset without relying on real-world motion capture, a finite element method-based simulator could be employed. Note that the markers are only used to create training and evaluation datasets, as well as evaluate the performance of DEFT by providing the ground truth. However, for real-world applications, the markers are not needed. For dual manipulation, a Franka Emika Research 3 robot and a Kinova Gen3 robot were used. An illustration of the experimental setup is shown in Figure 9. We used an OptiTrack motion capture system to record ground-truth vertex locations. Note in our implementation of DEFT we set $\epsilon=0.1$ (Theorem 5) and $\kappa=0.02$ (Algorithm 2). The parameters were selected to balance stability, convergence speed, and accuracy. The dataset is split with $\sim$75for training and $\sim$25for evaluation.